# Resonance Reduction Against Adversarial Attacks

### v1

@@ 11 $=([S]_{11}-[S]_{12}[S]_{22}^{-1}[S]_{21})^{-1}$ (14)
As a quick sanity check, note that when ${c}=0$ (which means that the main and the auxiliary networks are not connected), we have
@@ 11 $=(i\nu 2{\gamma}{\Omega}^{2}+{\Omega}^{2}-\nu^{2}I)^{-1}$ (15)
indicating that the steady-state response on the main network is equivalent to the one derived in equation (3), as expected. In Section VI we use this theoretical result to perform further numerical sanity check on the Auxiliary Graph Optimization objective function.
Although it is not possible to establish a configuration for the auxiliary graph that will guarantee it, we assume that ${\Omega}$ and ${\widetilde{\Omega}}$ are simultaneously diagonalizable. This, using (14), allows us to compute the eigenvalues of $[S^{-1}]_{11}$ as
@@
where ${\omega}_{k},{\widetilde{\omega}}_{k}$ are the $k$-th eigenvalues of ${\Omega}$ and ${\widetilde{\Omega}}$ respectively. Later on, we will show that there will be an approximation error between the computed expected value and the average squared $2$-norm of the steady-state response when ${\Omega}$ and ${\widetilde{\Omega}}$ are not simultaneously diagonalizable.
As before, we assume that $\mathbf{f}$ is being uniformly sampled from ($n-1$)-unit sphere and the adversarial agent only has imprecise information about the main graph as described in Section IV-B (i.e., it has no information about the auxiliary graph and hence the combined network). Using the results from the derivation of the Network Graph Optimization method (Section IV), we compute the expected-value of the random variable $\|{\mathbf{x}}_{s}\|^{2}_{2}$, hence the vulnerability of the main network when it is attached to an auxiliary network as:
[…]
where $\overline{s}_{k}$ denotes the complex conjugate of $s_{k}$ and $\rho_{\omega_{j}}(\nu)=\frac{h/\pi}{({\omega}_{j}-\nu)^{2}+h^{2}}$. A closed form expression for the integral is obtained using the Residue theorem with the same contour as before as described in Section IV-B. In order to use the Residue theorem as described, however, one needs to compute the roots of the quartic polynomial in $\nu$ in the denominator of the integrand and determine whether those roots have positive or negative imaginary parts. In this case a direct computation of that, even using a symbolic algebra toolbox, was not feasible because of the complexity of the problem. In order to simplify computation of the roots, we use linearization with respect to $\gamma$. The details of the computation are provided in Appendix -B. Corresponding calculations are performed using a symbolic mathematics toolbox. We omit the resulting expression for brevity.
@@
V-B Eigenspectrum Optimization of the Auxiliary Network Graph
@@ VI Results
  • Validate the accuracy of the objective functions, ${J}$ and ${\widetilde{J}}$, in representing the network vulnerability measured by $\mathbb{E}(\|{\mathbf{x}}_{s}\|^{2}_{2})$ for ${\mathbf{x}}_{s}$ defined on ${G}$ and ${\widetilde{G}}$, as described equations (9) and (17) respectively.
@@ VI-A Implementation Details and Setup
To solve the network graph and auxiliary graph eigenspectrum optimizations, we use the interior-point algorithm implemented within the fmincon function from MATLAB Optimization Toolbox [17].
@@ VI-A2 Adversarial Force Sampling
As described in Section IV-B, the forcing frequency needs to be sampled using a probability density function that is a uniformly weighted sum of multiple Cauchy distributions each of which are centered at the natural frequencies, $\{\omega_{j}\}_{j=1,\dots,n}$, with a constant spread of $h$. However, in order to perform this sampling, one needs to compute the inverse of the cumulative distribution function (c.d.f.) of the $\rho$ described in (5), which is computationally difficult. We thus adopt a practical method that is used to generate samples from mixture models as explained in [22]: Using an uniform probability of $1/n$ on all the natural frequencies, $\{\omega_{j}\}_{j=1,\dots,n}$, we first sample one of the natural frequencies, say $\omega_{s}$. Then the forcing frequency is sampled from a Cauchy distribution centred at $\omega_{s}$ and with a spread of $h$ using the inverse c.d.f. of Cauchy distribution, $\nu=\omega_{s}+h\tan(\pi(p-0.5)$, where $p$ is sampled from an uniform distribution over the unit interval $[0,1]$.
@@ VI-B1 Validation of the Objective Function
The objective function, ${J}$ (equation (9)), for the network graph eigenspectrum optimization problem is the expected value of the squared 2-norm of the steady-state response of the dynamic network subject to adversarial forcing with the stochastic model explained in Sections IV-AIV-B.
@@
Figure 8: The network graph eigenspectrum optimization problem is solved for instances featuring 100 RCGs, 100 RIGs and 100 social network graphs. On problems instances where the main network graphs are complete, the optimization consistently yielded larger relative decrease in the objective values.
Figure 9: Eigenvalue spectrum of a complete graph before and after optimization represented as histograms of the eigenvalues of the networks stiffness matrix, $L+\varepsilon I$. The objective value significantly decreased from 1.378 to 0.3778, yielding a 72.58% decrease as a result of the network graph optimization on a 100-node RCG. Qualitatively, as a result of the network graph optimization, the eigenvalue spectrum has become flatter.
@@
VI-B2 Parameter Analysis
Each eigenspectrum optimization problem on a network graph can be specified via a set of parameters regarding the second order dynamics of the network, the external forcing and the constraints. These parameters are the number of vertices on the graph ($n$), the number of edges on the graph ($n_{e}$), the minimum weight constraint ($w^{min}$), the stiffness constant ($\varepsilon$), the damping factor (${\gamma}$), and the spread of the external agent’s frequency distribution ($h$).
We analyze the effects of these parameters on the percentage reduction of objective value that can be achieved via the eigenspectrum optimization, hence the reduction in the vulnerability of the main network graph using the Network Graph Optimization method. For this purpose, we start with set of parameter values, $n=30$, $n_{e}=225$, $w^{min}=10^{-3}$, $\varepsilon=10$, ${\gamma}=10^{-6}$, $h=0.1$, and generate problem instances featuring both RCGs and RIGs where we vary one parameter and keep the rest constant. We solve for each problem instance and compute the percentage reductions in objective as %$\%d_{{J}}=\frac{|{J}^{0}-{J}^{*}|}{{J}^{0}}\times 100$ (where superscripts 00 and * denote initial and optimal objective values), which are plotted against the varying parameter values in Figure 7. Note that by comparing the percentage decrease in the objectives instead of the final objective values achieved, we are trying to isolate the effect of the parameters on the effectiveness of Network Graph Optimization method in reducing the vulnerability of a graph instead of trying to find the set of problem parameters that make the network the least vulnerable.
It can be observed from Figure 7 that a larger decrease in the objective value can be achieved as the number of vertices or the number of edges increase. Intuitively, more vertices and more edges correspond to more flexibility in distributing the weight resources, thus resulting in larger improvements to the vulnerability of the graph. As expected, larger stiffness yields better results, where the effect gets more significant with increased orders of magnitude. The spread of the external agent’s frequency distribution has a non-monotonic effect. As the spread gets smaller, the agent is able to pick the resonance frequencies more accurately, leaving the graph helpless against the attack, whereas a larger frequency spread corresponds to an agent that almost arbitrarily picks its frequencies, against which any modification of the graph based on reasoning would be less effective. Since the minimum weight constraint and the damping factor did not demonstrate a significant effect on the percentage decrease of the objective, corresponding plots are excluded. By observing the plots overall and the analysis on the number of edges, it is clear that the eigenspectrum optimization on a main network graph is more effective when the graph is complete. This behavior will become more apparent in the next section.
To demonstrate the overall effectiveness of eigenspectrum optimization on the main network graph in reducing the network vulnerability, we solve the optimization problem for RCGs, RIGs and Social graphs, and show that significant decrease in objective values can be achieved. We generate 100 RCGs and RIGs with $n$ sampled uniformly from the interval $[10,30]$ and $w_{p}$ sampled uniformly from the interval $[0.1,0.5]$. For the RIGs, we sampled $n_{e}$ from the interval $[n,n^{2}/4]$. The parameters associated with the network dynamics are the minimum weight, $w^{min}=0.001$, the stiffness constant, $\varepsilon=10$, the damping coefficient, ${\gamma}=10^{-6}$, and the adversarial agent’s frequency spread, $h=0.1$. The average percentage decrease in the objective value and the standard deviation across the problem instances featuring RCGs, RIGs and Social graphs are provided in Figure 8. Qualitatively, on the eigenspectrum of the graph, the optimization is manifested as a flattening of the eigenspectrum, as can be seen for the complete graph in Figure 10 and the Social graph in Figure 13.
@@ VI-B3 Demonstration of the Effectiveness of Network Graph Optimization
As mentioned before, network graph eigenspectrum optimization is more successful at reducing the objective value relative to the initial value of the objective when it is performed on complete graphs. A reason for this behavior is the greater vulnerability of the complete graphs to the resonance attacks, due to the fact that the natural frequencies of a complete graph are heavily accumulated around a value resulting in a peaky spectrum, compared to a relatively flatter/uniform distribution of the natural frequencies on an incomplete graph. A fewer number of optimization variables impose greater rigidity on incomplete graphs due to their fewer edges, whereas complete graphs, with their maximum possible number of edges, offer a greater flexibility in edge weight manipulations. Qualitatively, this is manifested by a lower relative flattening of the eigenspectrum in case of the incomplete Social graph (Figure 13) as compared to the complete graph (Figure 10).
@@ VI-B4 Numerical Second-Order Dynamics Simulation of the Main Network
We simulated (performed numerical integration of) the second-order dynamics on 100 different graphs, with both the initial and optimized weights for a complete graph and a social graph, and with varying forcing vectors and sampled forcing frequencies. The simulations were run until a steady state was achieved and the final steady-state amplitude was noted.
@@
Complete Graph Numerical Simulations: We considered an unoptimized complete graph with uniform edge weights with an added perturbation as detailed in Section  VI-A, as well as the corresponding optimized graph obtained using the network graph optimization method detailed in Section IV, and performed 100 numerical simulations on each of these graphs. The squared amplitude of $\mathbf{x}$ as a function of time for each of the 100 simulations, each normalized by the closed-form steady-state squared amplitude $\|\mathbf{x}_{s}\|^{2}$, is shown in Figure 10. Besides observing that the steady-state amplitudes of the numerical simulations match the computed closed-form values, we note that the unsteady amplitude in relation to the steady-state amplitude is noticeably lower in the optimized graph.
Figure 11: An embedding of the representative Facebook Social Subgraph with 173 vertices. The colors on the edges indicate weights after optimization.
Figure 12: Eigenvalue spectrum of the representative Facebook Social subgraph before and after optimization represented as histograms of the eigenvalues of the networks stiffness matrix, $L+\varepsilon I$. The objective value significantly decreased from 6.868 to 2.467 as a result of the network graph optimization on the 173-node Facebook social subgraph, corresponding to a 64.089% decrease in the objective value. As a result of the optimization, the eigenvalue spectrum has become smoother and flatter.
Figure 13: Normalized amplitude plot of 100 different simulations of the second-order dynamics for both the initial and optimized representative social graph, displaying their corresponding two-norm squared amplitude values over time, divided by the steady-state squared amplitude value calculated by the closed-form evaluation for the respective forcing vector and forcing frequency (i.e., $\frac{\|\mathbf{x}(t)\|^{2}}{\|\mathbf{x}_{s}\|^{2}}$). Since the two-norm squared amplitude value from the simulation of the second-order dynamics should converge to the same closed-form steady-state evaluation for the two-norm squared amplitude given the same corresponding forcing vector and forcing frequency, the values in the plot are expected to approach 1 indicated by the blue line. In these $100$ simulations, the initial (unoptimized) graph’s mean steady-state squared amplitude value is 3.768, while the optimized graph’s mean steady-state squared amplitude value is 1.922.
Social Graph Eigenspectrum Optimization: From the 100 social subgraphs generated, as described in Section VI-A1, a representative subgraph was chosen for numerical simulation. An embedding of the optimized graph is shown in Figure 11. The plot of the square amplitude as a function of time divided by the closed-form steady-state amplitude in Figure 13 reveal that numerical convergence was slower, with an average percent difference between the numerical steady-state amplitude and the closed form steady-state amplitude across the 100 samples being 0.0951% for the initial graph and 0.0433% for the optimized graph.
@@ VI-C1 Validation of the Objective Function
The objective function ${\widetilde{J}}$ for the auxiliary graph eigenspectrum optimization problem is the expected value of the squared 2-norm of the steady-state response corresponding to the main network vertices when the dynamic network is subject to stochastic adversarial forcing. To validate the accuracy of the objective function in representing the expected value, we generate $800$M adversarial forcing samples, evaluate the closed-form steady state responses for each sample using equation (13) and compute the average squared 2-norm of the responses. For the validation study, we generate a RCG with $n=10$, with $w_{p}=0.3$ and use it as the main network graph. The running average over the number of samples divided in multiple batches are provided in Figure 14.
@@
The problem instance generated for the validation study resulted in an optimized auxiliary graph for which ${\Omega}$ and ${\widetilde{\Omega}}^{*}$ are simultaneously diagonalizable. As a consequence, we observe that $\overline{\|{\mathbf{x}}_{s}\|^{2}_{2}}$ converge to the objective values for both the optimized and unoptimized combined networks. To demonstrate the fact that there will be an approximation error between $\overline{\|{\mathbf{x}}_{s}\|^{2}_{2}}$ and ${\widetilde{J}}$, when ${\Omega}$ and ${\widetilde{\Omega}}^{*}$ are not simultaneously diagonalizable, we perform another auxiliary graph eigenspectrum optimization based on a RIG with $n=10$, $n_{e}=25$ and $w_{p}=0.3$. The running average over the number of samples divided in multiple batches are provided in Figure 15.
@@
From Figures 14 and 15, it can be observed that over a large number of forcing samples, the average of the squared 2-norm of the steady-state responses is well approximated by the objective values when ${\Omega}$ and ${\widetilde{\Omega}}$ are simultaneously diagonalizable, whereas there exist an approximation error when these matrices are not simultaneously diagonalizable. Also, it can be seen that on the optimized graphs, the steady state responses have smaller amplitudes on average.

### v2

@@ 11 $=([S]_{11}-[S]_{12}[S]_{22}^{-1}[S]_{21})^{-1}$ (14)
(As a quick sanity check, note that when ${c}=0$, which means that the main and the auxiliary networks are not connected, we have
@@ 11 $=(i\nu 2{\gamma}{\Omega}^{2}+{\Omega}^{2}-\nu^{2}I)^{-1}$ (15)
indicating that the steady-state response on the main network is equivalent to the one derived in equation (3), as expected. In Section VI we use this theoretical result to perform further numerical sanity check on the Auxiliary Graph Optimization objective function.)
Since ${\Omega}$ and ${\widetilde{\Omega}}$ are simultaneously diagonalizable, using (14), allows us to compute the eigenvalues of $[S^{-1}]_{11}$ as
@@
According to the stochastic model explained in Section IV-A, $\mathbf{f}$ is being uniformly sampled from ($n-1$)-unit sphere and the adversarial agent only has imprecise information about the main graph (i.e., it has no information about the auxiliary graph and hence the combined network) leading to the probability density function (4) for the forcing frequency $\nu$.
Rest of the proof is similar to the proof of Proposition 1. We use Lemma 1 and 2, and equation (16) to compute the expected value with respect to $\mathbf{f}$. the result of the proposition then follows from the substitution of this expected value together with the p.d.f. (4) into equation (5). ∎
Later on, we will show that there will be an approximation error between the computed expected value and the average squared $2$-norm of the steady-state response when ${\Omega}$ and ${\widetilde{\Omega}}$ are not simultaneously diagonalizable.
A closed form expression for the integral in Proposition 3 is obtained using the Residue theorem with the same contour as before as described in Section IV-B. In order to use the Residue theorem as described, however, one needs to compute the roots of the quartic polynomial in $\nu$ in the denominator of the integrand and determine whether those roots have positive or negative imaginary parts. In this case a direct computation of that, even using a symbolic algebra toolbox, was not feasible because of the complexity of the problem. In order to simplify computation of the roots, we use linearization with respect to $\gamma$. The details of the computation are provided in Appendix -B. Corresponding calculations are performed using a symbolic mathematics toolbox. We omit the resulting expression for brevity.
@@
V-C Spectrum Optimization of the Auxiliary Network Graph
@@ VI Results
  • Validate the accuracy of the objective functions, ${J}$ and ${\widetilde{J}}$, in representing the network vulnerability measured by $\mathbb{E}(\|{\mathbf{x}}_{s}\|^{2}_{2})$ for ${\mathbf{x}}_{s}$ defined on ${G}$ and ${\widetilde{G}}$, as described in Proposition 1 and 3 respectively.
@@
  • Apply the network graph optimization to the communication network among a team of mobile robots between which the signal strength decays with increasing distance.
@@ VI-A Implementation Details and Setup
We solve the network graph and auxiliary graph spectrum optimizations using the interior-point algorithm [23].
@@ VI-A2 Adversarial Force Sampling
As described in Section IV-A, the forcing frequency needs to be sampled using a probability density function that is a uniformly weighted sum of multiple Cauchy distributions each of which are centered at the natural frequencies, $\{\omega_{j}\}_{j=1,\dots,n}$, with a constant spread of $h$. However, in order to perform this sampling, one needs to compute the inverse of the cumulative distribution function (c.d.f.) of the $\rho$ described in equation (4), which is computationally difficult. We thus adopt a practical method that is used to generate samples from mixture models as explained in [28]: Using an uniform probability of $1/n$ on all the natural frequencies, $\{\omega_{j}\}_{j=1,\dots,n}$, we first sample one of the natural frequencies, say $\omega_{s}$. Then the forcing frequency is sampled from a Cauchy distribution centred at $\omega_{s}$ and with a spread of $h$ using the inverse c.d.f. of Cauchy distribution, $\nu=\omega_{s}+h\tan(\pi(p-0.5)$, where $p$ is sampled from an uniform distribution over the unit interval $[0,1]$.
@@ VI-B1 Validation of the Objective Function
The objective function, ${J}$ (Proposition 1), for the network graph spectrum optimization problem is the expected value of the squared 2-norm of the steady-state response of the dynamic network subject to adversarial forcing with the stochastic model explained in Section IV-A.
@@
Figure 8: The network graph spectrum optimization problem is solved for instances featuring 100 RCGs, 100 RIGs and 100 social network graphs. On problems instances where the main network graphs are complete, the optimization consistently yielded larger relative decrease in the objective values.
VI-B2 Parameter Analysis
Each spectrum optimization problem on a network graph can be specified via a set of parameters regarding the second order dynamics of the network, the external forcing and the constraints. These parameters are the number of vertices on the graph ($n$), the number of edges on the graph ($n_{e}$), the minimum weight constraint ($w^{min}$), the stiffness constant ($\varepsilon$), the damping factor (${\gamma}$), and the spread of the external agent’s frequency distribution ($h$).
We analyze the effects of these parameters on the percentage reduction of objective value that can be achieved via the spectrum optimization, hence the reduction in the vulnerability of the main network graph using the Network Graph Optimization method. For this purpose, we start with set of parameter values, $n=30$, $n_{e}=225$, $w^{min}=10^{-3}$, $\varepsilon=10$, ${\gamma}=10^{-6}$, $h=0.1$, and generate problem instances featuring both RCGs and RIGs where we vary one parameter and keep the rest constant. We solve for each problem instance and compute the percentage reductions in objective as %$\%d_{{J}}=\frac{|{J}^{0}-{J}^{*}|}{{J}^{0}}\times 100$ (where superscripts 00 and * denote initial and optimal objective values), which are plotted against the varying parameter values in Figure 7. Note that by comparing the percentage decrease in the objectives instead of the final objective values achieved, we are trying to isolate the effect of the parameters on the effectiveness of Network Graph Optimization method in reducing the vulnerability of a graph instead of trying to find the set of problem parameters that make the network the least vulnerable.
It can be observed from Figure 7 that a larger decrease in the objective value can be achieved as the number of vertices or the number of edges increase. Intuitively, more vertices and more edges correspond to more flexibility in distributing the weight resources, thus resulting in larger improvements to the vulnerability of the graph. As expected, larger stiffness yields better results, where the effect gets more significant with increased orders of magnitude. The spread of the external agent’s frequency distribution has a non-monotonic effect. As the spread gets smaller, the agent is able to pick the resonance frequencies more accurately, leaving the graph helpless against the attack, whereas a larger frequency spread corresponds to an agent that almost arbitrarily picks its frequencies, against which any modification of the graph based on reasoning would be less effective. Since the minimum weight constraint and the damping factor did not demonstrate a significant effect on the percentage decrease of the objective, corresponding plots are excluded. By observing the plots overall and the analysis on the number of edges, it is clear that the spectrum optimization on a main network graph is more effective when the graph is complete. This behavior will become more apparent in the next section.
To demonstrate the overall effectiveness of spectrum optimization on the main network graph in reducing the network vulnerability, we solve the optimization problem for RCGs, RIGs and Social graphs, and show that significant decrease in objective values can be achieved. We generate 100 RCGs and RIGs with $n$ sampled uniformly from the interval $[10,30]$ and $w_{p}$ sampled uniformly from the interval $[0.1,0.5]$. For the RIGs, we sampled $n_{e}$ from the interval $[n,n^{2}/4]$. The parameters associated with the network dynamics are the minimum weight, $w^{min}=0.001$, the stiffness constant, $\varepsilon=10$, the damping coefficient, ${\gamma}=10^{-6}$, and the adversarial agent’s frequency spread, $h=0.1$. The average percentage decrease in the objective value and the standard deviation across the problem instances featuring RCGs, RIGs and Social graphs are provided in Figure 8. Qualitatively, on the spectrum of the graph, the optimization is manifested as a flattening of the spectrum, as can be seen for the complete graph in Figure 17 and the Social graph in Figure 17.
@@ VI-B3 Demonstration of the Effectiveness of Network Graph Optimization
As mentioned before, network graph spectrum optimization is more successful at reducing the objective value relative to the initial value of the objective when it is performed on complete graphs. A reason for this behavior is the greater vulnerability of the complete graphs to the resonance attacks, due to the fact that the natural frequencies of a complete graph are heavily accumulated around a value resulting in a peaky spectrum, compared to a relatively flatter/uniform distribution of the natural frequencies on an incomplete graph. A fewer number of optimization variables impose greater rigidity on incomplete graphs due to their fewer edges, whereas complete graphs, with their maximum possible number of edges, offer a greater flexibility in edge weight manipulations. Qualitatively, this is manifested by a lower relative flattening of the spectrum in case of the incomplete Social graph (Figure 17) as compared to the complete graph (Figure 17). An embedding of an optimized Social graph is shown in Figure 18
@@ VI-B4 Numerical Second-Order Dynamics Simulation of the Main Network
We simulated (performed numerical integration of) the second-order dynamics on 100 different graphs, with both the initial and optimized weights with varying forcing vectors and sampled forcing frequencies. The simulations were run until a steady state was achieved and the final steady-state amplitude was noted.
@@
Complete Graph Numerical Simulations: We considered an unoptimized complete graph with uniform edge weights with an added perturbation as detailed in Section  VI-A, as well as the corresponding optimized graph obtained using the network graph optimization method detailed in Section IV, and performed 100 numerical simulations on each of these graphs. The squared amplitude of $\mathbf{x}$ as a function of time for each of the 100 simulations, each normalized by the closed-form steady-state squared amplitude $\|\mathbf{x}_{s}\|^{2}$, is shown in Figure 9. Besides observing that the steady-state amplitudes of the numerical simulations match the computed closed-form values, we note that the unsteady amplitude in relation to the steady-state amplitude has less variation in the optimized graph.
VI-B5 Application of Network Graph Optimization to a Robot Network
We consider a team of $n$ mobile robots and their communication network described by a complete graph. The signal strength between robot $i$ and $j$ (represented by the edge weight $w_{ij}$) is computed as: $w_{ij}=\frac{A_{dist}}{\|r_{i}-r_{j}\|+\varepsilon_{dist}}$, where $A_{dist}$ and $\varepsilon_{dist}$ are some constants, and $r_{i}$, $r_{j}$ indicate the positions of robots $i$ and $j$.
We solve the optimization problem defined in Section IV-C and the constraints on the edge weights defined therein. In addition, we consider a physical constraint that prevents robot collision given as:
$\|r_{i}-r_{j}\|\geq d_{min}\quad\forall i,j\in\mathbb{Z}\text{ and }1\leq i,j\leq n$
Then, the goal is to optimally relocate each robot such that the objective value is minimized.
We consider three types of initial configurations for the robots: arbitrary placement within some bounding box, on a uniform grid, on a line. We generate 10 instances for each initial condition where some small random perturbation is applied to the robot locations. Following parameters are used for the experiments: $n=30$, $w^{min}=0.001$, $\varepsilon=1$, ${\gamma}=10^{-6}$, $h=0.1$, $A_{dist}=1$, $\varepsilon_{dist}=0.1$, $d_{min}=1$.
The mean and the standard deviation of the objective reduction achieved from each type of initial configuration is reported in Table I. Initial and optimal robot locations for one instance of the problem are provided in Figure 10.
TABLE I: Robot Network Optimization Results
Arbitrary Grid Line
Mean %$\%d_{{J}}$ $27.38$ $39.20$ $27.02$
Std. %$\%d_{{J}}$ $4.05$ $0.91$ $0.64$
Figure 10: Robots are initially arranged on a $6\times 5$ grid with some small perturbations. After the optimization, robots are relocated to minimize the objective.
@@ VI-C1 Validation of the Objective Function
The objective function ${\widetilde{J}}$ for the auxiliary graph spectrum optimization problem is the expected value of the squared 2-norm of the steady-state response corresponding to the main network vertices when the dynamic network is subject to stochastic adversarial forcing. To validate the accuracy of the objective function in representing the expected value, we generate $800$M adversarial forcing samples, evaluate the closed-form steady state responses for each sample using equation (13) and compute the average squared 2-norm of the responses. For the validation study, we generate a RCG with $n=10$, with $w_{p}=0.3$ and use it as the main network graph. The running average over the number of samples divided in multiple batches are provided in Figure 11.
@@
The problem instance generated for the validation study resulted in an optimized auxiliary graph for which ${\Omega}$ and ${\widetilde{\Omega}}^{*}$ are simultaneously diagonalizable. As a consequence, we observe that $\overline{\|{\mathbf{x}}_{s}\|^{2}_{2}}$ converge to the objective values for both the optimized and unoptimized combined networks. To demonstrate the fact that there will be an approximation error between $\overline{\|{\mathbf{x}}_{s}\|^{2}_{2}}$ and ${\widetilde{J}}$, when ${\Omega}$ and ${\widetilde{\Omega}}^{*}$ are not simultaneously diagonalizable, we perform another auxiliary graph spectrum optimization based on a RIG with $n=10$, $n_{e}=25$ and $w_{p}=0.3$. The running average over the number of samples divided in multiple batches are provided in Figure 12.
@@
From Figures 11 and 12, it can be observed that over a large number of forcing samples, the average of the squared 2-norm of the steady-state responses is well approximated by the objective values when ${\Omega}$ and ${\widetilde{\Omega}}$ are simultaneously diagonalizable, whereas there exist an approximation error when these matrices are not simultaneously diagonalizable. Also, it can be seen that on the optimized graphs, the steady state responses have smaller amplitudes on average.